OpenStudy (anonymous):

What is the similarity ratio of two cubes with volumes of 125 cubic meters and 216 cubic meters?

OpenStudy (anonymous):

like 1:5

OpenStudy (anonymous):

\[\sqrt[3]{\frac{125}{216}}=\frac{5}{6}\text{ }\text{or}\text{ 5:6 }\]